OpenStudy (anonymous):

Let the hcf & lcm of of two natural numbers a, b be 6 and 36 respectively and a + b = 42, then a - b is equal to

OpenStudy (anonymous):

it is hcf and lcm

OpenStudy (ikram002p):

hcf =gcd (greatest common devisor ) right ?

OpenStudy (ikram002p):

lcm(a,b)=a*b/gcd(a,b) 36=ab/6 ab=216

OpenStudy (ikram002p):

a=216/b a=42-b 216/b=42-b 216=42b-b^2 b^2-42+216=0 (solve for b)

OpenStudy (ikram002p):

b^2-42b+216=0 (solve for b)

OpenStudy (ikram002p):

(b-36)(b-6)=0 b=36 or b=6 case 1 b=36 a=216/36=6 a-b=-30 case 2 a=36 b=216/36=6 a-b=30

OpenStudy (anonymous):

@ganeshie

OpenStudy (ikram002p):

add 8 to tag him :P

OpenStudy (ikram002p):

like this @ganeshie8